# Multi-Domain Audio Question Answering Toward Acoustic Content Reasoning in The DCASE 2025 Challenge

### v1

@@ Abstract
We present Task 5 of the DCASE 2025 Challenge: an Audio Question Answering (AQA) benchmark spanning multiple domains of sound understanding. This task defines three QA subsets (Bioacoustics, Temporal Soundscapes, and Complex QA) to test audio-language models on interactive question-answering over diverse acoustic scenes. We describe the dataset composition (from marine mammal calls to soundscapes and complex real-world clips), the evaluation protocol (top-1 accuracy with answer-shuffling robustness), and baseline systems (Qwen2-Audio-7B, AudioFlamingo 2, Gemini-2-Flash). Preliminary results on the development set are compared, showing strong variation across models and subsets. This challenge aims to advance the audio understanding and reasoning capabilities of audio-language models toward human-level acuity, which are crucial for enabling AI agents to perceive and interact about the world effectively.
@@ 1 Introduction
Contemporary audio AI research increasingly aims to build systems capable of interactive audio understanding: models that not only recognize sound events but also reason about them [1]. The DCASE 2025 Challenge Audio Question Answering (AQA) task is motivated by this goal. It requires systems to interpret diverse acoustic scenes, incorporate relevant external knowledge such as facts or subjective acoustic information about sounds of marine mammals or everyday situations, and reason over both perceived (multiple) sound events and background information to generate appropriate answers.
@@
Figure 1: The proposed audio question answering (AQA) dataset with three categories: Bioacoustics QA (BQA), Temporal Soundscapes QA (TSQA), and Complex QA (CQA).
Compared to audio classification tasks [12, 13], which involve understanding audio content and assigning it to a predefined label, AQA requires additional capabilities, including the ability to interpret the question, incorporate relevant knowledge, and link it to the audio information to select the correct answer. Among prior DCASE tasks, automated audio captioning [14] is the most closely related to AQA. While both involve generating natural language from audio input, captioning typically focuses on high-level event descriptions. In contrast, AQA demands fine-grained reasoning, contextual interpretation, and instruction-following to answer specific questions grounded in both auditory signals and external knowledge. A comparison between these tasks is illustrated in Figure 2.
In this report, we present the details of the newly proposed DCASE 2025 Challenge AQA task. Specifically, we introduce the AQA dataset, describe the evaluation methodology, and present baseline model results for the task.
Figure 2: Causal graphs illustrating key differences among three audio understanding tasks. Shaded (gray) nodes represent observable variables, while unshaded nodes represent latent or implicit factors. (a) In Audio Classification, models map audio recordings to predefined labels based solely on surface acoustic features. (b) Automated Audio Captioning introduces an intermediate representation involving interactive events and conceptual understanding to generate natural language descriptions. (c) In Audio Question Answering (AQA), reasoning depends not only on the audio and the question but also on latent confounding factors such as audio information, interactive events, and external knowledge. These factors enrich the reasoning process required to derive accurate answers.
@@ 2 Dataset Details
The AQA task dataset for the DCASE 2025 Challenge comprises three distinct QA subsets, each following a multiple-choice question format. Most of the questions present four answer choices (while there are some two and three answer choices in Part 3), with exactly one correct option. Each subset is designed to assess a different aspect of audio understanding and reasoning. All subsets include a training set and a development (validation) set for model training and evaluation.
@@ 2.3 Part 3: Complex QA (CQA)
2.4 Dataset and Codebase Availability
All subsets are provided through a HuggingFace repository, [^1] https://huggingface.co/datasets/PeacefulData/2025_DCASE_AudioQA_Official which contains JSON-formatted QA files along with links to the corresponding audio clips. Audio for Part 1 is downloaded from the official Watkins Marine Mammal Sound Database using the provided script. For the other subsets, participants may alternatively download all audio files directly using the links provided in the HuggingFace repository. We also provide a Qwen2-Audio-7B AQA baseline for the DCASE 2025 challenge [^2] https://huggingface.co/PeacefulData/2025_DCASE_AudioQA_Baselines with an average AQA accuracy of 45.045.045.045.0%, which will be introduced in section 4.1.
@@ 3.1 Evaluation Metric
Submissions will be evaluated on a held-out evaluation set, which will be released on June 1, 2025. The primary evaluation metric is top-1 accuracy, defined as the proportion of questions for which the model’s predicted answer matches the ground truth. Leaderboard rankings will be determined by top-1 accuracy averaged across the three domains (i.e., the mean of the per-domain accuracies) as shown in the row 2 of Tab 1. If several entries achieve the same score within their 85% binomial confidence intervals, the tie will be broken using their sample-weighted accuracies as shown in the row 3 of Tab 1.
To assess robustness, we additionally evaluate each model by shuffling the order of answer choices multiple times. The best and worst accuracies across these permutations are reported to indicate the model’s sensitivity to the ordering of options.
3.2 Submission Guidelines
Each team may submit up to four systems. Each system must produce a .csv output file containing two columns: question (the question ID) and answer (the selected option). An accompanying .yaml file describing system metadata must also be provided for each submission. The final submission package should be a single .zip file that includes all system outputs, metadata files, and the team’s technical report. Care must be taken to ensure that all output files strictly follow the required format.
4 Baseline Systems and Results
We evaluated three representative audio-language models as baselines in a zero-shot setting on the development sets.
@@ 4.1 Baseline Systems
AudioFlamingo 2. AudioFlamingo 2 [7] is a 3-billion-parameter audio-centric large language model built upon a Flamingo-style cross-attention architecture. It integrates an enhanced CLAP [24] audio encoder with a language decoder. The model is trained using a multi-stage curriculum that incorporates both real and synthetic data. A key component of this training is the AudioSkills dataset, a large-scale synthetic corpus designed for audio-based instruction following and question answering, covering a broad range of reasoning skills.
@@ 4.3 Qualitative Analysis
Table 2: Hallucination examples in Qwen2-Audio-7B responses
@@
Which species is identified from the sound recording? A C Identified as a combination of three types of whales, but it is only one type of “Narwhal”
@@
In the first example, the model identifies sounds such as “mechanical fan” and “ticking clock” between 2.0s and 4.0s, none of which are part of the annotated ground truth or aligned with the provided answer choices. This suggests the model is either over-relying on statistical priors or misinterpreting background textures as discrete sound events. This indicates a common issue in models lacking temporal resolution calibration. The second case demonstrates Multi-Selection, where the model incorrectly identifies the sounds of “Narwhal” as a combination of three types of whales. However, there is only one correct answer. In the third case, although the model selects the correct answer (“C”), the justification involves an event (“laughter”) that is not among the choices.
Table 3: Correct prediction in Qwen2-Audio-7B responses
@@
Which species and vocalization type can be identified? A A States presence of Pantropical Spotted Dolphin (correct target)
Which species is identified from the recording? A A Lists Grampus griseus, matching correct answer
Correct Predictions (Win). The model accurately identifies fine-grained acoustic patterns and species-specific vocalizations. Notably, it demonstrates precise spectral reasoning (e.g., identifying modulated high-frequency components) and species recognition in bioacoustic contexts. These successes indicate the model’s potential for perceptual grounding in structured audio tasks. Examples are presented in Table 3.
Table 4: Wrong prediction in Qwen2-Audio-7B responses
Question Correct Model Model Explanation (Excerpt)
Which species and vocalization type is most consistent with the sound recording? A C Describes Sperm Whale clicks, contradicting correct species and vocalization type
@@
What type of vocalization is most prominently identified? A C Identifies “rapidly modulated clicking” at 8.67–9.02s, differing from the correct “pulsed squeal”
What sound occurs immediately after the keys dropping? B A Mentions “telephone bell ringing” rather than “cough”
Incorrect Predictions (Wrong). Incorrect predictions arise when the model either misattributes acoustic sources or fails to temporally align events with their corresponding semantic categories. As shown in Table 4, these errors span species misclassification, timestamp imprecision, and confusion between perceptually similar vocalizations. Such failures underscore limitations in fine-grained auditory discrimination and temporal grounding.
@@ 5 Conclusion
In this report, we introduce the AQA task for the DCASE 2025 Challenge, which requires systems to answer multiple-choice questions based on audio inputs. The task comprises three subsets, Bioacoustics QA, Temporal Soundscapes QA, and Complex QA, each targeting different aspects of audio understanding and reasoning. We outline the evaluation methodology, with top-1 accuracy as the primary metric and permutation-based accuracy for robustness. We also present three zero-shot baseline models: Qwen2-Audio-7B, AudioFlamingo 2, and Gemini-2.0-Flash. Results show overall low performance and substantial variation across models and subsets, highlighting the challenge’s difficulty and the need for more advanced approaches.
We encourage participants to explore a range of techniques, including efficient fine-tuning methods and inference-time strategies (e.g., task-activating prompting [26, 27]). We hope this challenge fosters progress toward more robust, generalizable, and human-like audio understanding and reasoning.
Acknowledgements
We thank the maintainers of the underlying datasets that made this work possible. Specifically, the Bioacoustics QA subset utilizes recordings from the Watkins Marine Mammal Sound Database (Woods Hole Oceanographic Institution) [15], and we gratefully acknowledge the permission granted to use the database for research and challenge purposes. The Temporal Soundscapes QA subset incorporates sounds from the NIGENS General Sound Events Dataset [16], the L3DAS23 Challenge Set [17], and the TAU 2019 Spatial Sound Events dataset [18], all of which are publicly available. The Complex QA subset draws audio from AudioSet [19] and MiraData [20]. We sincerely acknowledge the contributors of these resources.

### v2

# Multi-Domain Audio Question Answering Benchmark Toward Acoustic Content Reasoning

@@ Abstract
We introduce a multi-domain Audio Question Answering (MD-Audio) benchmark designed to advance research in multi-domain sound understanding. The task is organized into three subsets (Bioacoustics, Temporal Soundscapes, and Complex QA), each targeting distinct aspects of interactive question answering over diverse acoustic environments. The benchmark encompasses a broad range of auditory material, including marine mammal vocalizations, environmental soundscapes, and complex real-world audio recordings. We further detail the evaluation protocol, which employs top-1 accuracy alongside an answer-shuffling robustness criterion, and provide baseline results using state-of-the-art audio-language models (Qwen2-Audio-7B, AudioFlamingo 2, and Gemini-2-Flash). The non-archrival benchmark is released as an open-source resource and used for the DCASE 2025 AQA Challenge. Preliminary analyses on the development set reveal substantial variability in model performance across both tasks and domains. Ultimately, this benchmark111huggingface.co/datasets/PeacefulData/2025_DCASE_AudioQA_Official is intended to foster progress in the auditory comprehension and reasoning abilities of audio-language models, capabilities that are critical for enabling systems to perceive, interpret, and interact with the acoustic world in a manner approaching human-level acuity.
@@ 1 Introduction
Contemporary audio AI research increasingly aims to build systems capable of interactive audio understanding: models that not only recognize sound events but also reason about them [1]. The MD-Audio is motivated by this goal. It requires systems to interpret diverse acoustic scenes, incorporate relevant external knowledge such as facts or subjective acoustic information about sounds of marine mammals or everyday situations, and reason over both perceived (multiple) sound events and background information to generate appropriate answers.
Fig. 1: The proposed multi-domain audio question answering (MD-Audio) dataset with three categories: Bioacoustics QA, Temporal Soundscapes QA, and Complex QA.
@@
Compared to audio classification tasks [12, 13], which involve understanding audio content and assigning it to a predefined label, AQA requires additional capabilities, including the ability to interpret the question, incorporate relevant knowledge, and link it to the audio information to select the correct answer. A comparison between these tasks is illustrated in Figure 2. In this study, we present the details of the newly proposed MD-Audio and describe the evaluation methodology, and present baseline model results for the task.
Fig. 2: Causal graphs illustrating key differences among three audio understanding tasks. Shaded (gray) nodes represent observable variables, while unshaded nodes represent latent or implicit factors. (a) In Audio Classification, models map audio recordings to predefined labels based solely on surface acoustic features. (b) Automated Audio Captioning introduces an intermediate representation involving interactive events and conceptual understanding to generate natural language descriptions. (c) In Audio Question Answering (AQA), reasoning depends not only on the audio and the question but on latent confounding factors such as audio information, interactive events, and external knowledge. These factors enrich the reasoning process to derive accurate answers.
@@ 2 Dataset Details
The AQA task dataset  [^2] https://huggingface.co/PeacefulData/2025_DCASE_AudioQA_Baselines comprises three distinct QA subsets, each following a multiple-choice question format. Most of the questions present four answer choices (while there are some two and three answer choices in Part 3), with exactly one correct option. Each subset is designed to assess a different aspect of audio understanding and reasoning. All subsets include a training set and a development (validation) set for model training and evaluation.
@@ 3.1 Evaluation Metric
Submissions will be evaluated on a held-out evaluation set, which will be released on June 1, 2025. The primary evaluation metric is top-1 accuracy, defined as the proportion of questions for which the model’s predicted answer matches the ground truth. Leaderboard rankings will be determined by top-1 accuracy averaged across the three domains (i.e., the mean of the per-domain accuracies) as shown in the row 2 of Tab 1. If several entries achieve the same score within their 85% binomial confidence intervals, the tie will be broken using their sample-weighted accuracies as shown in the row 3 of Tab 1. From the participants’ result, we noticed inference-time strategies (e.g., task-activating prompting [20, 21]) and policy based post-training adaptation would make model attained competative performance.
@@ 3.2 Baseline Systems
AudioFlamingo 2. AudioFlamingo 2 [7] is a 3-billion-parameter audio-centric large language model built upon a Flamingo-style cross-attention architecture. It integrates an enhanced CLAP [25] audio encoder with a language decoder. The model is trained using a multi-stage curriculum that incorporates both real and synthetic data. A key component of this training is the AudioSkills dataset, a large-scale synthetic corpus designed for audio-based instruction following and question answering, covering a few reasoning skills.
@@
Table 1: Zero-shot accuracy (% with $\pm$ 85% confidence interval) of baselines on the dev-set across the subsets of the AQA task.
@@ 3.4 Qualitative Analysis
Table 2: Hallucination responses in Qwen2-Audio-7B
@@
In the first example, the model identifies sounds such as “mechanical fan” and “ticking clock” between 2.0s and 4.0s, none of which are part of the annotated ground truth or aligned with the provided answer choices. This suggests the model is either over-relying on statistical priors or misinterpreting background textures as discrete sound events. This indicates a common issue in models lacking temporal resolution calibration. In the following case, although the model selects the correct answer (“C”), the justification involves an event (“laughter”) that is not among the choices.
Table 3: Correct & wrong responses in Qwen2-Audio-7B
@@
Correct and Wrong Predictions . The model accurately identifies fine-grained acoustic patterns and species-specific vocalizations. Notably, it demonstrates precise spectral reasoning (e.g., identifying modulated high-frequency components). Incorrect predictions arise when the model either misattributes acoustic sources or fails to temporally align events with their corresponding semantic categories.
@@ 4 Conclusion
We introduce the AQA task for the DCASE 2025 Challenge, which requires systems to answer multiple-choice questions based on audio inputs. The task comprises three subsets, Bioacoustics QA, Temporal Soundscapes QA, and Complex QA, each targeting different aspects of audio understanding and reasoning. We outline top-1 accuracy for ranking and permutation-based accuracy for robustness.